OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

@Zale101 this is my last problem would u please help

OpenStudy (anonymous):

i haven't attempted it yet but could u walk me through it?@Zale101 @Nnesha

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@some.random.cool.kid

OpenStudy (some.random.cool.kid):

you called :}

OpenStudy (anonymous):

yes i need help with this someone please helpppp

OpenStudy (some.random.cool.kid):

ok

Nnesha (nnesha):

perpendicular line have opposite slope flip and change sign what is your slope in the equation that they gave you ??

OpenStudy (some.random.cool.kid):

ok same method zale tought you

OpenStudy (anonymous):

this is the whole equation

OpenStudy (some.random.cool.kid):

plug in the numbers like you did before to start.

OpenStudy (anonymous):

which slope the y-y1=m(x-x1)

OpenStudy (anonymous):

im lost what nnesha said

Nnesha (nnesha):

you have slope intercept form equation not point  what is slope here ???

Nnesha (nnesha):

y = mx + b here m is slope always with x

OpenStudy (anonymous):

wait thats the answer? y= -4x-2?

OpenStudy (some.random.cool.kid):

yeah also to add wenever asked which is the slope look for the number with the x just in case you didnt know.

Nnesha (nnesha):

this is equation that you have in the question

OpenStudy (anonymous):

how did you get to the answer can u please draw out the steps?

OpenStudy (anonymous):

lol didnt see that

OpenStudy (anonymous):

can someone please draw out the steps I'm kinda lost and seeing it would help

Nnesha (nnesha):

yes what is slope in that equation we need that ??